obtain the dimension of k in the equation W=1\2Kx^2 Here w is work x is change in lenghth  dimension is  and how

Grade:11

1 Answers

Prathik
4 years ago
Given equation   W=1/2kx2
Where W------) Work
             x-------)length
wwrite the dimensional formula of work
i.e W=Fxs=(M1L1T-2)(L)1=M1L2T-2
Now we know
                        W=1/2kx2
                     M1L3T-2=k(L)2   (as 1/2 is
dimensionless)
                       So  k=M1 L1 T-2